A can do a piece of work in two-thirds as many day as B,and B can do it in three-fourths as many days as C. Working together, they can do it in 2 and ten-thirteenths days. In how many days can each of them alone do the work.
Answer by nyc_function(2741)

2 10/13 = 36/13
C takes n days to do one job.
B takes (3/4) n days to do one job.
A takes (2/3) [(3/4) n] days to do one job. [= n/2 days to do one job.]
C can complete 1/n jobs in a day.
B can complete 4/(3n) jobs in a day.
A can complete 2/n jobs in a day.
36/13 days * [1/n jobs/day + 4/(3n) jobs/day + 2/n jobs/day] = 1 job.
solve for n, then plug-n-chug into:
C takes n days to do one job.
B takes (3/4) n days to do one job.
A takes n/2 days to do one job.